OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3) solve and show work

OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3) distributive property, or get rid of the parenthesis: -2x + 6 >= 5 - x - 3 combine like terms: -2x + 6 >= 2 - x add x to both sides: -2x + 6 + x > = 2 - x + x -x + 6 >= 2 subtract 6 from both sides: -x + 6 - 6 >= 2 - 6 -x >= -4 multiply by -1 and switch the direction of the inequality: x <= 4